Hi i need to show steps on how to do this problem and the problem is 4x^2-3x+10=4. Can you show me step by step and explain clearly in detail to show all steps. thanks
Answer by lynnlo(4176)

4x^2-3x+10=4
MOVE -4 TO THE LEFT
4X^2-3X+10-4=0
ADD FRACTIONS OR INTEGERS TOGETHER
4X^2-3X+6=0